# Improving Cramér–Rao Bound And Its Variants: An Extrinsic Geometry Perspective

### v1

@@ 2 Preliminaries
In this section, we fix notation and the geometric framework, including square-root embedding, connections, and second fundamental form, used in the sequel. The section is concluded with a brief sketch of our plan to use these ideas to improve upon the CRB and some higher-order versions of the kind worked out in [8, 11].
2.1 Notation, square-root embedding, and assumptions
@@
exist and are in $L^{2}(\mu)$. Further, these are assumed to be linearly independent. Differentiation under the integral sign is permitted; with $T(\cdot):\mathbb{E}_{\theta}[T]<\infty$,
@@
The Fisher–Rao metric on the 1-dimensional parameter manifold is $g(\partial_{\theta},\partial_{\theta})=\mathcal{I}(\theta)$, with $\partial_{\theta}$ denoting the coordinate vector field. Equip $s(\Theta)$ with the induced metric from $L^{2}(\mu)$. The pullback metric from the square-root embedding then equals $\tfrac{1}{4}\mathcal{I}(\theta)$.
@@ 2.2 Geometric preliminaries
Note that we work with the fixed Hilbert space $L^{2}(\mu)$ because it provides a fixed inner product and a metric-compatible Levi–Civita connection. The ambient connection in $L^{2}(P_{\theta})$ is not compatible with the metric there. Of course, one could enforce metric-compatibility by hand (by “redefining the Christoffel symbols”), but then we are no longer working with the Levi–Civita connection of the ambient Hilbert space – the geometry is compromised – and we would be simply importing back the $L^{2}(\mu)$ structure in disguise.
@@
Define the “second fundamental vector" by taking the normal component of the ambient derivative:
$\mathrm{\Pi}_{m}\equiv\mathrm{II}(\eta_{1},\eta_{m}):=\eta_{m+1}-\sum_{k=1}^{m}\Gamma^{\mathrm{proj}}_{k}\,\eta_{k}.$ (6)
For $m=1$, this is just the standard second fundamental form. By construction, $\mathrm{II}(\eta_{1},\eta_{m})\in\mathcal{T}_{m}(\theta)^{\perp}$, i.e.
$\langle\mathrm{II}(\eta_{1},\eta_{m}),\eta_{i}\rangle=0\qquad\text{for }1\leq i\leq m.$
@@ 2.3 Faà di Bruno and exponential Bell polynomials
Recall $s_{\theta}(x)=\sqrt{f(x;\theta)}$. For brevity, we suppress the dependence on $x$ and $\theta$ when no ambiguity arises. Set
@@
Application to the square-root embedding.
@@ 2.4 Outlook and road ahead
Consider any unbiased estimator $T(X)$ ($X\sim P_{\theta}$) of the scalar parameter $\theta$ and define the centered estimation error $Z_{0}:=T(X)-\theta$ viewed as an element of $L^{2}(P_{\theta})$ or, equivalently, focus on the vector $\widetilde{Z}_{0}:=Z_{0}s_{\theta}\in L^{2}(\mu)$. Similarly, also define $\widetilde{Y}_{1}:=Y_{1}s_{\theta},\widetilde{Y}_{2}:=Y_{2}s_{\theta},\dots,\widetilde{Y}_{m}:=Y_{m}s_{\theta}$ and $\widetilde{\mathcal{T}}_{m}:=\mathrm{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{m}\}.$
@@
Then the squared norm $\|\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\|^{2}$ is the projection-based portion of the variance (for $m=1$ it reproduces the CRB, as just mentioned); the residual $\|R_{m}\|^{2}$ is the unexplained variance by the $m$-term projection. Projecting $R_{m}$ further onto the curvature span $\mathcal{N}_{m}$ (see (7)) yields an explicit curvature correction. In summary, we may better the CRB by considering the curvature-induced residual for $m=1$. We emphasize again that this idea is distinct from the method of Bhattacharyya [8] which uses higher-order derivatives of the score and does not exploit the extrinsic geometry, as we do, via the square root embedding.
With $m>1$, Bhattacharyya’s [8, 9, 10]/Pillai and Sinha’s result [11] can be considered as improving the CRB using the subspace $\widetilde{\mathcal{T}}_{m}$, and not $\mathcal{T}_{m}=\mathrm{span}\{\eta_{1},\dots,\eta_{m}\}$, in the projection above (strictly speaking, Bhattacharyya/Pillai–Sinha consider the ratios of the derivatives of the density $f(x;\theta)$ to the density function, and not directly the derivatives of the log-likelihood as already mentioned). In any case, we prefer to work with $\mathcal{T}_{m}$ in developing our preliminary curvature-correction result (cf. Theorem 5 below) for reasons related to those quoted in Remark 1; in short, this approach offers an intuitive and correct geometric picture. One may ask: How does the subspace $\mathcal{T}_{m}$ compare against $\widetilde{\mathcal{T}}_{m}$ or, equivalently, how does our theorem relate to those of the nature of Bhattacharyya/Pillai–Sinha? While the spaces $\widetilde{\mathcal{T}}_{1}=\mathcal{T}_{1}$, the answer is a bit more subtle in the $m>1$ case, and we refer interested readers to Section 4.3 for some stronger (in terms of improvement in variance lower bounds) results that shed light on this scenario and Section 4.4 for further discussion. Our strongest result for $m>1$, which betters bounds of the kind obtained by Bhattacharyya/Pillai and Sinha, appears in Proposition 8.
@@ 3 Curvature-corrected variance bounds – CRB case
The proof of the $m>1$ case in Theorem 5 essentially covers that of $m=1$ as well, but we state the $m=1$ case separately since it corresponds to improving the classical CRB.
Theorem 2 (Residual variance and CRB improvement, $m=1$).
@@
4 Higher-order CRB and curvature corrections in the square-root embedding
Needless to (re)state, we keep Assumption 1 so that all quantities below are well defined in the fixed Hilbert space $L^{2}(\mu)$. We first state the Hilbert-space formulation of the Bhattacharyya bound in $L^{2}(P_{\theta})$ as conceptualized in [11] with the caveat that we use derivatives of the log likelihood and not the ratios $\frac{\partial_{\theta}^{k}f}{f}$, show its equivalent formulation in the ambient space $L^{2}(\mu)$, and then present a basic curvature-corrected bound formulated relative to the jet subspace $\mathcal{T}_{m}=\operatorname{span}\{\eta_{1},\dots,\eta_{m}\}$; the core idea here is the same as for the $m=1$ case. This bound is finally strengthened in Proposition 8 in Section 4.3, where it is explicitly shown that we improve over the Bhattacharyya-type bounds in the framework of Pillai and Sinha using derivatives of the raw score.
4.1 Bhattacharyya/Pillai–Sinha-type bounds in $L^{2}(P_{\theta})$ and its square-root representation
The following theorem formalizes, strengthens, and rephrases the core insight of Pillai and Sinha [11], who construct a sequence of variance lower bounds using higher-order score functions systematically covering the Bhattacharyya variants [8]. We claim no real novelty in the following result, but include the same anyway for the following reasons: While their work asserts that equality in the $m$-th-order bound results if estimator error lies in the span of score functions up to order $m$, our formulation provides a characterization of equivalence using Hilbert space projections, and sets the stage for further improvements to be derived later via basic Riemannian geometric arguments.
@@ 4.1 Bhattacharyya/Pillai–Sinha-type bounds in $L^{2}(P_{\theta})$ and its square-root representation
Theorem 3 (Hilbert space characterization of higher-order CRB [8, 11]).
@@
where the $Y_{i}$ are assumed to be linearly independent. Let $G^{Y}\in\mathbb{R}^{m\times m}$ be the Gram matrix with entries $G^{Y}_{ij}:=\langle Y_{i},Y_{j}\rangle_{\theta}$, and define $b\in\mathbb{R}^{m}$ by $b_{i}:=\langle Z_{0},Y_{i}\rangle_{\theta}$. Denote the variance $\sigma^{2}:=\|Z_{0}\|_{\theta}^{2}=\operatorname{Var}_{\theta}[T(X)]$, and define the matrix:
@@
    $B^{(m)}$ is singular; specifically, $\left(\left[(B^{(m)})^{-1}\right]_{11}\right)^{-1}=0$.
@@
(ii) $B^{(m)}$ is singular and $\left(\left[(B^{(m)})^{-1}\right]_{11}\right)^{-1}=0$.
@@
It is obvious, again for $m=1$, that the general bound $b^{\top}(G^{Y})^{-1}b$ becomes:
@@
Lemma 4 (Square-root isometry and score-image representation).
@@
4.2 Curvature-corrected variance bound, $m>1$
We now state and prove the curvature-corrected bound formulated relative to the jet subspace $\mathcal{T}_{m}$, and not $\widetilde{\mathcal{T}}_{m}$, since this really is the natural generalization of the $m=1$ case (see also Remark 1). Thus we work in the fixed Hilbert space $L^{2}(\mu)$ and the jet subspace $\mathcal{T}_{m}$. The geometric refinement we use is again similar in spirit to the $m=1$ scenario: project the estimator residual onto the jet space, then project the residual onto the curvature span inside the normal bundle.
@@ 4.2 Curvature-corrected variance bound, $m>1$
$\mathrm{II}_{m}:=\mathrm{II}(\eta_{1},\eta_{m})=\eta_{m+1}-\sum_{k=1}^{m}\Gamma^{\mathrm{proj}}_{k}\,\eta_{k},$
@@
$\operatorname{Var}_{\theta}[T]\geq\mathcal{B}^{(m)}_{\mathrm{jet}}+\|\operatorname{Proj}_{\mathcal{N}_{m}}(\widetilde{Z}_{0})\|^{2}=:\mathcal{C}^{(m)},$ (16)
@@
$\widetilde{\mathcal{T}}_{k}=\operatorname{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{k}\},\qquad\mathcal{T}_{k}=\operatorname{span}\{\eta_{1},\dots,\eta_{k}\}.$
The Bhattacharyya/Pillai–Sinha-type projection bounds correspond to projecting the estimator residual onto $\widetilde{\mathcal{T}}_{k}$ (equivalently, projecting $Z_{0}$ onto $\mathcal{T}_{k}^{Y}$ in $L^{2}(P_{\theta})$). That projection yields $\mathcal{B}^{(k)}_{\mathrm{score}}=\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{k}}\widetilde{Z}_{0}\|^{2}$.
The curvature correction requires projecting onto the jet subspace $\mathcal{T}_{k}=\operatorname{span}\{\eta_{1},\dots,\eta_{k}\}$, which is the natural subspace to be considered with the induced connection and second fundamental vectors. Even though $\mathcal{T}_{k},\,\widetilde{\mathcal{T}}_{k}$ are both $k$-dimensional subspaces under assumptions of linear independence, in general $\mathcal{T}_{k}\neq\widetilde{\mathcal{T}}_{k}$ for $k\geq 2$, so $\mathcal{B}^{(k)}_{\mathrm{jet}}$ and $\mathcal{B}^{(k)}_{\mathrm{score}}$ are different quantities. Only when the linear spans exactly coincide do the two formulations agree. The one-dimensional spaces $\operatorname{span}\{\widetilde{Y}_{1}\}$ and $\operatorname{span}\{\eta_{1}\}$ coincide because $\eta_{1}=\tfrac{1}{2}\widetilde{Y}_{1}$, which is why the classical CRB is unaffected by the choice. Nevertheless, it is obvious that our result in Theorem 5 offers curvature-aware improvement over $\mathcal{B}^{(m)}_{\mathrm{score}}$ when $\mathcal{B}^{(m)}_{\mathrm{jet}}=\mathcal{B}^{(m)}_{\mathrm{score}}$ with $m>1$. What can we say otherwise? Some answers are given in the next section, where we argue that it is not (very) beneficial to only consider the single normal vector in $\mathrm{span}\{\Pi_{m}\}$.
4.3 Improving over Bhattacharyya/Pillai–Sinha-type bounds $\mathcal{B}^{(m)}_{\mathrm{score}}$
We start by noting the following, most of which we already met earlier.

### v2

@@ 2 Preliminaries
In this section, we fix notation and the geometric framework, including square root embedding, connections, and second fundamental form, used in the sequel. The section is concluded with a brief sketch of our plan to use these ideas to refine the CRB and some higher-order versions of the kind worked out in [8].
2.1 Notation, square root embedding, and assumptions
@@
exist in the Fréchet sense in $L^{2}(\mu)$. Further, these are assumed to be linearly independent. Differentiation under the integral sign is permitted; with $T(\cdot):\mathbb{E}_{\theta}[T]<\infty$,
@@
The Fisher–Rao metric on the 1-dimensional parameter manifold is $g(\partial_{\theta},\partial_{\theta})=\mathcal{I}(\theta)$, with $\partial_{\theta}$ denoting the coordinate vector field. Equip $s(\Theta)$ with the induced metric from $L^{2}(\mu)$. The pullback metric from the square root embedding then equals $\tfrac{1}{4}\mathcal{I}(\theta)$.
@@ 2.2 Geometric preliminaries
Note that we work with the fixed Hilbert space $L^{2}(\mu)$ because it provides a fixed inner product and a metric-compatible Levi–Civita connection. The ambient connection in $L^{2}(P_{\theta})$ is not compatible with the metric there. Of course, one could enforce metric-compatibility by hand (by “redefining the Christoffel symbols”), but then we are no longer working with the Levi–Civita connection of the ambient Hilbert space – the geometry is compromised – and we would be simply importing back the $L^{2}(\mu)$ structure in disguise. The key role played by the square root map is also highlighted in our work [38] in the context of the associated jet bundle geometry.
@@
Define the “second fundamental vector” by taking the normal component of the ambient derivative:
$\mathrm{\Pi}_{m}:=\eta_{m+1}-\sum_{k=1}^{m}\Gamma^{\mathrm{proj}}_{k}\,\eta_{k}.$ (6)
For $m=1$, this is just the standard second fundamental form $\mathrm{II}(\eta_{1},\eta_{1})$. By construction, $\mathrm{II}_{m}\in\mathcal{T}_{m}(\theta)^{\perp}$, i.e.
$\langle\mathrm{II}_{m},\eta_{i}\rangle=0\qquad\text{for }1\leq i\leq m.$
@@
Note that $\Pi_{m}$ may be interpreted as the remnant of the “acceleration” $\eta_{m+1}$ after subtracting out the “tangential” part in $\mathcal{T}_{m}$. A more formal presentation of the second fundamental vector may be found in our follow up [38], where we cast it within the Cartan prolongation framework applied to the hierarchy of jet bundles.
@@ 2.3 Faà di Bruno and exponential Bell polynomials
Recall $s_{\theta}(x)=\sqrt{f(x;\theta)}$. For brevity, we suppress the dependence on $x$ and $\theta$ when there is no ambiguity. Set
@@
Application to the square root embedding.
@@ 2.4 Outlook and road ahead
Consider any unbiased estimator $T(X)$, where $X\sim P_{\theta}$, of the scalar parameter $\theta$ and define the centered estimation error $Z_{0}:=T(X)-\theta$ viewed as an element of $L^{2}(P_{\theta})$ or, equivalently, focus on the vector $\widetilde{Z}_{0}:=Z_{0}s_{\theta}\in L^{2}(\mu)$. Similarly, also define $\widetilde{Y}_{1}:=Y_{1}s_{\theta},\widetilde{Y}_{2}:=Y_{2}s_{\theta},\dots,\widetilde{Y}_{m}:=Y_{m}s_{\theta}$ and $\widetilde{\mathcal{T}}_{m}:=\mathrm{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{m}\}.$
@@
Then the squared norm $\|\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\|^{2}$ is the projection-based portion of the variance (for $m=1$ it reproduces the CRB, as just mentioned); the residual $\|R_{m}\|^{2}$ is the unexplained variance by the $m$-term projection. Projecting $R_{m}$ further onto the curvature span $\mathcal{N}_{m}$ (see (7)) yields an explicit curvature correction. In summary, we may refine the CRB by considering the curvature-induced residual for $m=1$. We emphasize again that this idea is distinct from the method of Bhattacharyya [8] which uses higher-order derivatives of the likelihood and does not exploit the extrinsic geometry, as we do, via the square root embedding. Those works that adopt an extrinsic differential geometric viewpoint, such as those of Efron [17] (see also Theorem 4.4 in [16]), are statements about asymptotic variance bounds.
With $m>1$, Bhattacharyya-type results in the literature [8, 9, 10] can be considered as improving the CRB using the subspace $\widetilde{\mathcal{T}}_{m}$, and not $\mathcal{T}_{m}=\mathrm{span}\{\eta_{1},\dots,\eta_{m}\}$, in the projection above. In any case, we prefer to work with $\mathcal{T}_{m}$ in developing our Bhattacharyya-type bound, which can be viewed as a curvature-correction theorem (cf. Theorem 5 below), for reasons related to those quoted in Remark 1; in short, this approach offers an intuitive and correct geometric picture. One may ask: How does the subspace $\mathcal{T}_{m}$ compare with $\widetilde{\mathcal{T}}_{m}$ or, equivalently, how does our Bhattacharyya-type theorem relate to those that appeared before, using log-likelihood derivatives? While the spaces $\widetilde{\mathcal{T}}_{1}=\mathcal{T}_{1}$, the answer is a bit more subtle in the $m>1$ case, and we refer interested readers to Section 4.3 for some results that shed light on this scenario and Section 4.4 for further discussion. Our conclusion, which improves bounds of the nature of Bhattacharyya using log-likelihood, appears in Proposition 8 where we specifically use the relationship between the $\eta_{j}$ and $Y_{j}$ given in (10).
@@ 3 Curvature-corrected variance bounds – CRB case
The proof of the $m>1$ case in Theorem 5 essentially covers that of $m=1$ as well, but we state the $m=1$ case separately since it corresponds to addressing inefficiency with respect to the classical CRB.
Theorem 2 (Residual variance and CRB refinement, $m=1$).
@@
4 Higher-order CRB and curvature corrections in the square root embedding
Needless to (re)state, we keep Assumption 1 so that all quantities below are well defined in the fixed Hilbert space $L^{2}(\mu)$. We first state the Hilbert-space formulation of the Bhattacharyya bound in $L^{2}(P_{\theta})$ (as conceptualized in other works such as [12], for example) with the caveat that we use derivatives of the log-likelihood and not the ratios $\frac{\partial_{\theta}^{k}f}{f}$, show its equivalent formulation in the ambient space $L^{2}(\mu)$, and then present a basic curvature-corrected bound formulated relative to the jet subspace $\mathcal{T}_{m}=\operatorname{span}\{\eta_{1},\dots,\eta_{m}\}$; the core idea here is the same as for the $m=1$ case. Taking into account the normal directions in $\mathcal{T}_{m}^{\perp}$ other than $\Pi_{m}$ (which uses $\eta_{m+1}$) in Proposition 8, we explicitly show that improvements over Bhattacharyya-type bounds using derivatives of the raw score of the same order $m$ can be achieved by comparing the subspaces $\mathcal{T}_{m}$ and $\widetilde{\mathcal{T}}_{m}$.
4.1 Bhattacharyya-type bounds in $L^{2}(P_{\theta})$ and its square root representation
The following theorem rephrases the core insight of other authors (for instance, [12, 13, 15]), who construct a sequence of variance lower bounds using higher-order score functions systematically covering the Bhattacharyya variants [8]. We claim no real novelty in the following result, but include the same anyway for the following reasons: While their work asserts that equality in the $m$-th-order bound results if estimator error lies in the span of score functions up to order $m$, our formulation provides a characterization of equivalence using Hilbert space projections, and sets the stage for further improvements to be derived later via basic Riemannian geometric arguments.
@@ 4.1 Bhattacharyya-type bounds in $L^{2}(P_{\theta})$ and its square root representation
Theorem 3 (Hilbert space characterization of higher-order CRB [8]).
@@
where the $Y_{i}$ are assumed to be linearly independent and $m\geq 1$. Let $G^{Y}\in\mathbb{R}^{m\times m}$ be the Gram matrix with entries $G^{Y}_{ij}:=\langle Y_{i},Y_{j}\rangle_{\theta}$, and define $b\in\mathbb{R}^{m}$ by $b_{i}:=\langle Z_{0},Y_{i}\rangle_{\theta}$. Denote the variance $\sigma^{2}:=\|Z_{0}\|_{\theta}^{2}=\operatorname{Var}_{\theta}[T(X)]$, and define the matrix:
@@
    $B^{(m)}$ is singular; specifically, $\left(\left[(B^{(m)})^{-1}\right]_{11}\right)^{-1}=0$ (under the convention that $\tfrac{1}{\infty}=0$).
@@
(ii) $B^{(m)}$ is singular and $\left(\left[(B^{(m)})^{-1}\right]_{11}\right)^{-1}=0$ under the convention $\frac{1}{\infty}=0$.
@@
For $m=1$, it is obvious that the general bound $b^{\top}(G^{Y})^{-1}b$ becomes:
@@
Lemma 4 (square root isometry and score-image representation).
@@
4.2 Curvature-corrected Bhattacharrya-type bound, $m>1$
We now state and prove the curvature-corrected bound, essentially our Bhattacharyya-type bound, formulated relative to the jet subspace $\mathcal{T}_{m}$, and not $\widetilde{\mathcal{T}}_{m}$, since this really is the natural generalization of the $m=1$ case (see also Remark 1). Thus we work in the fixed Hilbert space $L^{2}(\mu)$ and the jet subspace $\mathcal{T}_{m}$. The geometric refinement we use is again similar in spirit to the $m=1$ scenario: project the estimator residual onto the jet space, then project the residual onto the curvature span inside the normal bundle.
@@ 4.2 Curvature-corrected Bhattacharrya-type bound, $m>1$
$\mathrm{II}_{m}:=\eta_{m+1}-\sum_{k=1}^{m}\Gamma^{\mathrm{proj}}_{k}\,\eta_{k},$
@@
$\operatorname{Var}_{\theta}[T]\geq\mathcal{B}^{(m)}_{\mathrm{jet}}+\|\operatorname{Proj}_{\mathcal{N}_{m}}(\widetilde{Z}_{0})\|^{2}=:\mathcal{C}^{(m)}=\mathcal{B}^{(m+1)}_{\mathrm{jet}}$, (16)
@@
$\widetilde{\mathcal{T}}_{k}=\operatorname{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{k}\},\qquad\mathcal{T}_{k}=\operatorname{span}\{\eta_{1},\dots,\eta_{k}\},$
which are not the same for $k>1$. The Bhattacharyya-type projection bounds usually considered in the literature correspond to projecting the estimator residual onto $\widetilde{\mathcal{T}}_{k}$ (equivalently, projecting $Z_{0}$ onto $\mathcal{T}_{k}^{Y}$ in $L^{2}(P_{\theta})$). That projection yields $\mathcal{B}^{(k)}_{\mathrm{score}}=\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{k}}\widetilde{Z}_{0}\|^{2}$.
Interpretation of incorporating higher-order derivatives in a Bhattacharyya-type bound as curvature corrections requires projecting onto the jet subspace $\mathcal{T}_{k}=\operatorname{span}\{\eta_{1},\dots,\eta_{k}\}$, which is the natural subspace to be considered with the induced connection and second fundamental vectors. Even though $\mathcal{T}_{k},\,\widetilde{\mathcal{T}}_{k}$ are both $k$-dimensional subspaces under assumptions of linear independence, in general $\mathcal{T}_{k}\neq\widetilde{\mathcal{T}}_{k}$ for $k\geq 2$, so $\mathcal{B}^{(k)}_{\mathrm{jet}}$ and $\mathcal{B}^{(k)}_{\mathrm{score}}$ are different quantities. Only when the linear spans exactly coincide do the two formulations agree. The one-dimensional spaces $\operatorname{span}\{\widetilde{Y}_{1}\}$ and $\operatorname{span}\{\eta_{1}\}$ coincide because $\eta_{1}=\tfrac{1}{2}\widetilde{Y}_{1}$, which is why the classical CRB is unaffected by the choice. Nevertheless, it is obvious that our result in Theorem 5 offers a curvature-aware refinement of $\mathcal{B}^{(m)}_{\mathrm{score}}$ when $\mathcal{B}^{(m)}_{\mathrm{jet}}=\mathcal{B}^{(m)}_{\mathrm{score}}$ with $m>1$. Can we compare them in the general scenario where $\mathcal{T}_{k}\neq\widetilde{\mathcal{T}}_{k}$ for $k\geq 2$? This is what we seek to answer in the next section.
4.3 Improving over Bhattacharyya-type bounds $\mathcal{B}^{(m)}_{\mathrm{score}}$
What we proved in Theorem 5 is also a Bhattacharyya-type bound with the jets $\eta_{k}$, which was interesting for the reason that, in the sequence of bounds $\mathcal{B}_{\mathrm{jet}}^{(1)}\leq\mathcal{B}_{\mathrm{jet}}^{(2)}\leq\mathcal{B}_{\mathrm{jet}}^{(3)}\leq\dots$ ($\mathcal{B}_{\mathrm{jet}}^{(1)}$ equalling the CRB), we get
$\mathcal{B}_{\mathrm{jet}}^{(m+1)}-\mathcal{B}_{\mathrm{jet}}^{(m)}=\|\operatorname{Proj}_{\mathcal{N}_{m}}(\widetilde{Z}_{0})\|^{2},$
where the right side is precisely the curvature correction beyond order $m$. An interesting question is: Staying with derivatives of order $m$ and not considering the next higher derivative, can we demonstrate improvement over Bhattacharyya-type bounds that use the log-likelihood derivatives, i.e. $\mathcal{B}^{(m)}_{\mathrm{score}}$? For this purpose, we argue that it is beneficial to consider the structural relationship between the $\eta_{j}$ and $\widetilde{Y}_{j}$, subsequently incorporating normal directions in $\mathcal{T}_{m}^{\perp}$ other than the single normal vector in $\mathrm{span}\{\Pi_{m}\}$, which uses the $(m+1)$-th order derivative. That is, our goal here is ultimately to compare the subspaces $\mathcal{T}_{m}$ and $\widetilde{\mathcal{T}}_{m}$ for $m>1$.
We start by noting the following, most of which we met earlier.